Calculate the equilibrium income level Y given the following information:

Autonomous Consumption = 174

Marginal Propensity to Save = 0.48

Investment = 155


1
Expert's answer
2021-09-17T12:05:49-0400

C=174, I=155

"Y = C + I \\

Y =174 + 155 = 329"